The number of bacteria growing in an incubation culture increases with time according to B(t) = 3300(2)x, where x is time in days. Find the number of bacteria when x = 0 and x = 3.
Can you please help solve?

Answer by stanbon(75887)

The number of bacteria growing in an incubation culture increases with time according to B(t) = 3300(2)^x, where x is time in days. Find the number of bacteria when x = 0 and x = 3.
---
B(0) = 3300*2^0 = 3300*1 = 3300
------------------------------------
B(3) = 3300*2^3 = 3300*8 = 26,400
==================
Cheers,
Stan H.